OpenStudy (anonymous):

sqrt t (sqrt t - 6) = -9

OpenStudy (atlas):

let sqrt(t) =x solve for x now and then find t from x

OpenStudy (atlas):

\[x(x-6) = -9; where x =\sqrt t\] Hope you can do the rest

jhonyy9 (jhonyy9):

continue ... x(x-6)=-9 x^2 -6x = -9 x^2 -6x +9 =0 (x-3)(x-3)=0 (x-3)^2 =0 x-3 = 0 x = 3

jhonyy9 (jhonyy9):

so than continue x= sqrt(t) squared both sides you get x^2 = t 3^2 = t t = 9